2. The score on the entrance test for a well known law school has a mean score of 200 points and a standard deviation of 50 points. At value should the lowest passing score be set if the school wishes only 2.5% of those taking the entrance test to pass?

Answer by stanbon(48568)

The score on the entrance test for a well known law school has a mean score of 200 points and a standard deviation of 50 points. At what value should the lowest passing score be set if the school wishes only 2.5% of those taking the entrance test to pass?
--------------
1st: Find the z-score on the left tail of the normal curve that has
2.5% of the area to its left. That score is -1.96.
---------
2nd: Using the form z = (x-u)/s, solve for "x":
-1.96 = (x-200)/50
x-200 = -98
x = 102
------------
Conclusion: The lowest passing score should be 102
=============================
Cheers,
Stan H.